In the economy of Keynesian Island, autonomous consumption expenditure is $50 million, and the marginal propensity to consume is 0.8. Investment is $160 million, government expenditure is $190 million, and net taxes are $250 million. Investment, government purchases, and taxes are constant−they do not vary with income. The island does not trade with the rest of the world.

a. Draw the aggregate expenditure curve. (7 marks)

b. What is equilibrium real GDP for Keynesian Island? (3 marks)

c. What is the size of the multiplier in Keynesian Island's economy? (2 marks)

d. If the government increases its purchases by $200 million, what will be the change in the economy's equilibrium real GDP? (3 marks)
1
Expert's answer
2018-10-23T11:15:09-0400

C0 = $50 million, c = 0.8, I = $160 million, G = $190 million, T = $250 million.

a. In the aggregate expenditure model, equilibrium is the point where the aggregate supply and aggregate expenditure curve intersect.

b. Equilibrium real GDP for Keynesian Island is:

AE = C + I + G = Y, C = C0 + c*(Y - T),

Y = 50 + 0.8*(Y - 250) + 160 + 190,

0.2Y = 200,

Y = $1000 million.

c. The size of the multiplier in Keynesian Island's economy is:

m = 1/(1 - c) = 1/(1 - 0.8) = 5.

d. If the government increases its purchases by $200 million, then the change in the economy's equilibrium real GDP is 200*5 = $1000 million.